OpenStudy (anonymous):

what would be my ordered pairs when graphing a parabola of y=(x+3)^2+1

OpenStudy (shamil98):

y = a(x-h)^2 + k where (h,k) = vertex you have y = 1(x-(-3))^2 + 1 so (-3,1) is your vertex

OpenStudy (anonymous):

ok thank you and I have tried for x: -1, -2, -3, -4, -5 and if I were to continue to -6 the ordered pair is too high for my graph, also if I try 0 the ordered pairs is again to large for my graph.

OpenStudy (shamil98):

when x = 0 y = (0+3)^2 + 1 y = 10 (0,10) just make your graph represent larger numbers per dash or something

OpenStudy (anonymous):

the graph is automatically designed appropriately for the answers of its questions therefore I am unable to go higher than its limits and it can not be modified. my graph goes from y -axis: 9 to -9 and x-axis: 9 to -9

OpenStudy (shamil98):

not much i can do then.

OpenStudy (anonymous):

ok thank you anyways

OpenStudy (luigi0210):

Is this regular algebra or is this conics?